OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP: Determining the molar mass of a gas, with M=mRT/PV. m=.947 g, T=102 degrees C, and V= 212.85 mL. P=30.09 inHg. How do I find M? I keep getting a wierd decimal as a mol. How would I find the molar mass? Please help. I'm very stuck.

OpenStudy (anonymous):

m=.947 g T=102 C=375 K V=212.85 mL=0.21285 L

OpenStudy (anonymous):

well what R do u use for that?

OpenStudy (anonymous):

.0821?

OpenStudy (anonymous):

0.0821 what? unit?

OpenStudy (anonymous):

.08206 Latm/molK

OpenStudy (anonymous):

right so we need to change P=30.09 inHg.

OpenStudy (anonymous):

Into atm? Is it necessary to change units?

OpenStudy (anonymous):

yes absolutely...! because look at ur R units...what do u see?

OpenStudy (anonymous):

Oh...atm!

OpenStudy (anonymous):

yes so we need to change units

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

\[30.09 \ \text{inHg}\times \frac{25.4 \ \text{mmHg}}{1\ \text{inHg}} \times \frac{\text{1 atm}}{\text{760 mmHg}}=1.005\ \text{atm}\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Okay, so now how do I calculate M

OpenStudy (anonymous):

ok \[M=\frac{mRT}{PV}=\frac{0.947\times0.08206\times375}{1.005\times 0.21285}\]

OpenStudy (anonymous):

What do you get as the final answer? (So I can compare...)

OpenStudy (anonymous):

136 g/mol

OpenStudy (anonymous):

Okay, that's what I got. Thanks for the help!

OpenStudy (anonymous):

np :)